OpenStudy (anonymous):

Use completing the square to solve the equation x2 - 10x = -8.Which equation is used in the process? x2- 10x + 25 = 17 x2- 10x + 25 = -8 x2- 10x - 25 = -33

OpenStudy (campbell_st):

well the 1st step is to add half the coefficient of x then square it to both sides so the coefficient is -10 -10/2 = -5 and(-5)^2 = 25 so add 25 to both sides of the equation then write the equation as (x -5)^2 = 17 take the square root of both sides then add 5 so \[x = 5 + \sqrt{17}\]